OpenStudy (anonymous):

Mona needs the shingles for her roof to be right triangles . the lengths of the sides of the shingles are 15, 20, and 25. which equation can mona use to determine if the shingles are right triangles?

OpenStudy (anonymous):

\( 15^2+20^2=25^2 \) if that returns true, they're right triangles, otherwise they're not

OpenStudy (anonymous):

ty so if i have a choice to choose the best answer (15 =20)2 =25 2

OpenStudy (anonymous):

wrote that incorrectly sorry i meant to write (15+20)2 =25second power

OpenStudy (anonymous):

that's not the same

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

so that is incorrect

OpenStudy (anonymous):

yes, sorry